In the binomial replacement branching model with P(S) = q + ps , let  T = inf{n: Zn=0.

(a) Show that P[T=n] for n≥1 is  p^{n-1} \cdot q.

(b) Find P[T=n] for  0 < ? = 7.
Answer #1

In the binomial replacement branching model with P(S) = q + ps , let  T = inf{n: Zn=0.

(a) Show that P[T=n] for n≥1 is  b -ud.

(b) Find P[T=n] for  0 < ? = 7

In the binomial seplacement branching model with PO =+ps, let T = inf{n: 2:03. a). Grow that p{tin] for nslis phiq. Ans PS) =2. Find P[ian] assuming Z. =iso, Aly P[T =n91 20 =10) =P[inf Enizn =0} =n/zosiso] =P[2n 50;Zm1 =2, ..., Z=1, Z=1, Zori] = P(2